OpenStudy (anonymous):

What is the missing number? –3.4 + (–8 + 2) = [ + (–8)] + 2

OpenStudy (anonymous):

@Abhisar

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

(-12)+(-6)=x-8+2 -18=x-6 Try the rest please

OpenStudy (anonymous):

so what is the missing number

OpenStudy (anonymous):

hello @gaos

OpenStudy (anonymous):

X=-12